# FedHK-MVFC: Federated Heat Kernel Multi-View Clustering

### v1

@@ 4.5 Optimization Framework
  • Boundedness and Continuity: The objective function $J_{FedHK-MVFC}^{\ell}$ is continuous and bounded below. This follows from the fact that FKED values are bounded in $[0,1]$ by construction (Eq. 5), ensuring the weighted sum in the objective function remains finite. The exponential decay in the heat kernel guarantees that distances approach finite limits as data points move infinitely far apart.
  • Constraint Compactness: The constraint set is compact and convex. The membership matrix constraints $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$ with $\mu_{[\ell]ik}^{*}\in[0,1]$ define probability simplices, which are compact convex sets in $\mathbb{R}^{c(\ell)}$. Similarly, the view weight constraints $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$ with $v_{[\ell]h}\in[0,1]$ form a compact convex simplex in $\mathbb{R}^{s(\ell)}$. The cluster center parameters are unconstrained but remain bounded due to the data distribution.
  • Unique Subproblem Solutions: Each subproblem (updating $U$, AA, or $V$ while fixing others) has a unique solution. For membership updates, the strictly positive denominators in Eq. 28 ensure uniqueness since FKED values are always positive for distinct data points and cluster centers. For cluster center updates, the weighted exponential terms in Eq. 29 create a strictly convex weighted least-squares problem with a unique minimum. For view weight updates, the strictly decreasing nature of the power function with exponent $1/(\alpha-1)$ in Eq. 30 guarantees uniqueness of the normalized solution.
These conditions are satisfied by construction, as the FKED function is continuous and bounded, the probability simplex constraints define compact convex sets, and the update rules yield unique solutions when the denominators are non-zero (guaranteed by the strict positivity of FKED values and the non-degeneracy assumption that no data points are identical to cluster centers).
@@ 4.6 FedHK-MVFC Algorithms
The FedHK-MVFC algorithm is designed to operate in a federated learning environment, where each client processes its own multi-view data and shares only the necessary model parameters with the server. The algorithm consists of three main components: data preparation, initialization procedures, and the main federated learning process. The first component, data preparation, ensures that each client has a consistent and validated dataset. The second component initializes the global and local parameters, including view weights and cluster centers. The third component implements the main federated learning process, where clients iteratively update their local models based on the shared parameters from the server.
@@ 4.6.1 Data Preparation and Separation Sets
The preparation process encompasses systematic validation of multi-view datasets across participating clients. Each client undergoes rigorous dataset structure verification, confirming adequate sample sizes relative to clustering requirements and validating dimensional consistency across views. Cross-client feature alignment ensures compatibility between institutions while preserving the complementary nature of multi-view information.
Statistical quality assessment forms a critical component, involving outlier detection through Mahalanobis criteria and principled missing value imputation where data completeness exceeds acceptable thresholds. Feature-wise normalization with numerical stability safeguards maintains data integrity while enabling meaningful inter-client comparisons. The framework computes comprehensive quality metrics including global completeness scores and consistency measures across participating institutions.
Federation readiness certification provides formal validation that datasets meet the stringent requirements for collaborative analysis. The process generates essential federation metadata encompassing client counts, sample distributions, view configurations, and clustering parameters. Quality assurance protocols ensure that only datasets meeting predefined completeness and consistency thresholds proceed to the federated learning phase.
The separation sets emerge naturally from this preparation framework, where each client maintains its validated local dataset $X_{\left[\ell\right]}=\{x_{\left[\ell\right]ij}^{h}\}_{i=1}^{n(\ell)}$ across all available views. View-specific dimensionality $d_{\left[\ell\right]}^{h}$ and target cluster counts $c(\ell)$ are established during this phase, enabling seamless transition to initialization and iterative learning phases. Algorithm 2 provides the complete implementation framework for this comprehensive data preparation protocol.
@@ 4.6.2 Initialization Procedures
  • Initialize global view weights $V_{global}^{(0)}$ with equal distribution across views.
@@
For detailed implementation, refer to Algorithm 3. This algorithm outlines the server and client-side initialization procedures, ensuring that all clients are prepared for the federated learning process.
@@ 4.6.3 Main Federated Learning Process
The main federated learning process involves multiple communication rounds between the server and clients. It ensures that each client updates its local model based on the global model parameters received from the server. The clients perform local computations, including heat-kernel enhanced clustering, and send their model updates back to the server. The server aggregates these updates to refine the global model, which is then redistributed to all clients for the next round of learning. The main federated learning process is structured as follows:
@@
  • The process repeats for a predefined number of communication rounds or until convergence criteria are met.
In this framework, two servers are used to coordinate the federated learning process and manage the global model updates. The first server collects model updates from all clients, aggregates them, and updates the global cluster centers and view weights. The second server is responsible for distributing the updated global model parameters back to the clients for the next round of learning. The unified architecture of the two servers allows for efficient coordination and communication throughout the federated learning process. As a final step, the server evaluates the global model’s performance and adjusts the training strategy as needed. As it reaches convergence, the server outputs the final global model and personalized models for each client, along with their final membership matrices. And finally, the user can deploy a pattern of federated learning across different client devices into a production environment, ensuring that the models are trained and updated without compromising data privacy. A collective unifying framework is established to facilitate seamless integration and deployment of the federated learning models. This framework encompasses the entire lifecycle of federated learning, from model training to deployment, and provides the necessary tools and protocols for efficient collaboration among clients.
@@
  • Server-Side Components: The servers manage the global model, coordinate communication between clients, and handle aggregation of model updates.
@@ 4.6.4 Server Aggregation and Distribution
For the server able to efficiently aggregate and distribute model updates, the following steps are implemented:
@@
The aggregation process ensures that the global model is a representative update based on the contributions from all clients. This is crucial for maintaining the performance and accuracy of the federated learning system. For the server aggregation, we utilize a weighted averaging approach, where the contributions from each client are weighted based on their local data size or importance. The mathematical formulation for the weighted aggregation can be expressed as follows:
@@ 4.6.6 Statistics Aggregation
The statistical also includes the computation of weighted cluster centers and view weights for each client:
@@
These statistics provide insights into the local data distribution, cluster quality, and view importance for each client. They can be used to inform the server about the clients’ data characteristics and model performance. These statistics are shared with the server, which can use them to adjust aggregation weights, normalize global parameters, or detect data heterogeneity. For example, the server may assign higher aggregation weights to clients with lower variance, or use the global mean and variance to standardize cluster centers and view weights:
@@
where $\mu_{global}^{h}$ and $\sigma_{global}^{h}$ are the global mean and variance for view $h$. These global statistics can be used to normalize the local model updates, ensuring that the aggregated model is robust to variations in client data distributions. The server can then use these statistics to adjust the aggregation process, ensuring that the global model is robust and representative of the clients’ data. This statistical aggregation can also help in identifying outliers or anomalies in the client updates, allowing for more informed decision-making during the aggregation process.
@@ 4.6.8 The Computational Complexity of FedHK-MVFC
$O\left(\sum_{\ell=1}^{M}\left[n(\ell)^{2}d_{\left[\ell\right]}^{h}s(\ell)+n(\ell)c(\ell)s(\ell)+n(\ell)d_{\left[\ell\right]}^{h}c(\ell)s(\ell)\right]\right)$
@@ 5.1 Implementation of Privacy-Preserving Mechanisms
5.1.1 Data Locality and Federated Model Updates
In FedHK-MVFC, raw patient data never leaves the local hospital premises. Each hospital $\ell$ maintains its local multi-view dataset $X_{[\ell]}$ and only shares model parameters. The privacy-preserving update mechanism can be formalized as:
@@ 5.1.1 Data Locality and Federated Model Updates
where no raw data $x_{[\ell]ij}^{h}$ is transmitted. For example, Hospital A with ECG data, Hospital B with MRI scans, and Hospital C with genetic profiles only exchange aggregated cluster centers and view weights, not individual patient records.
5.1.2 Differential Privacy Integration
To provide formal privacy guarantees, FedHK-MVFC incorporates differential privacy by adding calibrated noise to model updates. The differentially private cluster center update becomes:
@@ 5.1.2 Differential Privacy Integration
5.1.3 Secure Aggregation Protocol
FedHK-MVFC employs a secure aggregation protocol where individual client contributions are cryptographically protected during the aggregation process. The server computes:
@@ 5.1.3 Secure Aggregation Protocol
where $\text{Encrypt}(\cdot)$ represents homomorphic encryption and $\text{SecAgg}(\cdot)$ denotes the secure aggregation function that allows computation on encrypted data without decryption.
5.1.4 Practical Privacy Example
Consider a concrete example where three hospitals cluster cardiovascular patients:
  • Hospital A: $n(1)=500$ patients with ECG data ($d_{[1]}^{1}=12$ leads)
  • Hospital B: $n(2)=300$ patients with MRI data ($d_{[2]}^{1}=1024$ voxel features)
  • Hospital C: $n(3)=400$ patients with genetic data ($d_{[3]}^{1}=50$ SNP markers)
Using FedHK-MVFC with $\epsilon=1.0$ differential privacy:
1. Each hospital computes local heat-kernel coefficients and cluster updates 2. Gaussian noise with variance $\sigma^{2}=\frac{2\log(1.25/10^{-5})}{1.0^{2}}\approx 23.03$ is added to updates 3. Encrypted model parameters are sent to the server 4. The server performs secure aggregation without accessing individual hospital data 5. Updated global parameters are redistributed for the next federated round
5.1.5 Privacy Budget Management
FedHK-MVFC implements adaptive privacy budget allocation across communication rounds. The total privacy budget $\epsilon_{total}$ is distributed as:
@@ 5.1.5 Privacy Budget Management
where $T$ is the total number of communication rounds. This allocation provides stronger privacy in early rounds when model updates are more informative, while maintaining convergence guarantees.
5.1.6 Communication Efficiency and Privacy Trade-offs
The privacy-preserving mechanisms in FedHK-MVFC achieve a favorable trade-off between privacy protection and communication efficiency. Compared to traditional federated approaches, FedHK-MVFC reduces communication overhead by 70% through:
@@ 5.1.6 Communication Efficiency and Privacy Trade-offs
The privacy-utility analysis shows that with $\epsilon=1.0$, clustering accuracy decreases by only 2-3% compared to the non-private version, while providing strong privacy guarantees for sensitive healthcare data.
@@ 6.1 Synthetic Multi-View Data Generation Framework
6.1.1 Mathematical Formulation of Synthetic Data Generation
Our synthetic data generator creates two-view, four-cluster datasets where each view exhibits distinct geometric patterns. This design philosophy ensures that the clustering algorithms must effectively integrate complementary information across views rather than relying on a single dominant view.
View 1: Parametric Geometric Structures
For View 1, we generate four clusters with distinct parametric shapes:

### v2

@@ 4.5 Optimization Framework
  • Boundedness and Continuity: The objective function $J_{FedHK-MVFC}^{\ell}$ is continuous and bounded below. This follows from the fact that FKED values are bounded in $[0,1]$ by construction (as defined in the federated KED formulation), ensuring that the weighted sum in the objective function remains finite. The exponential decay in the heat kernel guarantees that distances approach finite limits as data points move infinitely far apart. Furthermore, since all terms are non-negative and weighted by normalized coefficients, the objective function is bounded below by zero.
  • Constraint Compactness: The constraint set is compact and convex. The membership matrix constraints $\sum_{k=1}^{c(\ell)}\mu_{[\ell]ik}^{*}=1$ with $\mu_{[\ell]ik}^{*}\in[0,1]$ define probability simplices, which are compact convex sets in $\mathbb{R}^{c(\ell)}$. Similarly, the view weight constraints $\sum_{h=1}^{s(\ell)}v_{[\ell]h}=1$ with $v_{[\ell]h}\in[0,1]$ form a compact convex simplex in $\mathbb{R}^{s(\ell)}$. The cluster center parameters are unconstrained but remain bounded due to the data distribution and the heat kernel weighting, which assigns negligible influence to points far from cluster centers.
  • Unique Subproblem Solutions: Each subproblem (updating $U$, AA, or $V$ while fixing others) has a unique solution. For membership updates, the strictly positive denominators in Eq. 43 ensure uniqueness since FKED values are always positive for distinct data points and cluster centers (the exponential term never reaches exactly zero or one for finite distances). For cluster center updates, the weighted exponential terms in Eq. 44 create a strictly convex weighted least-squares problem with a unique minimum, as the composite weights $w_{[\ell]ijk}^{h}$ are strictly positive and the weighted squared error is a strictly convex function. For view weight updates, the strictly decreasing nature of the power function with exponent $1/(\alpha-1)$ in Eq. 45 guarantees uniqueness of the normalized solution, since the monotonic transformation preserves the ordering of view clustering qualities.
These conditions are satisfied by construction, as the FKED function is continuous and bounded, the probability simplex constraints define compact convex sets, and the update rules yield unique solutions when the denominators are non-zero (guaranteed by the strict positivity of FKED values and the non-degeneracy assumption that no data points are identical to cluster centers). The alternating optimization converges to a stationary point that satisfies the KKT conditions, representing a local optimum of the federated clustering objective. ∎
@@ 4.6 FedHK-MVFC Algorithms
The FedHK-MVFC algorithm is designed to operate in a federated learning environment, where each client processes its own multi-view data and shares only the necessary model parameters with the server. The algorithm consists of three main components: data preparation, initialization procedures, and the main federated learning process. The first component, data preparation, ensures that each client has a consistent and validated dataset. The second component initializes the global and local parameters, including view weights and cluster centers. The third component implements the main federated learning process, in which clients iteratively update their local models based on the parameters shared from the server.
@@ 4.6.1 Data Preparation and Separation Sets
The preparation process encompasses the systematic validation of multi-view datasets across participating clients. Each client undergoes rigorous dataset structure verification, confirming adequate sample sizes relative to clustering requirements and validating dimensional consistency across views. Cross-client feature alignment ensures compatibility between institutions while preserving the complementary nature of multi-view information.
Statistical quality assessment forms a critical component, involving outlier detection through Mahalanobis criteria and principled missing value imputation where data completeness exceeds acceptable thresholds. Feature-wise normalization with numerical stability safeguards maintains data integrity while enabling meaningful inter-client comparisons. The framework computes comprehensive quality metrics, including global completeness scores and consistency measures across participating institutions.
Federation readiness certification provides formal validation that datasets meet the stringent requirements for collaborative analysis. The process generates essential federation metadata, encompassing client counts, sample distributions, view configurations, and clustering parameters. Quality assurance protocols ensure that only datasets meeting predefined completeness and consistency thresholds proceed to the federated learning phase.
The separation sets emerge naturally from this preparation framework, where each client maintains its validated local dataset $X_{\left[\ell\right]}=\{x_{\left[\ell\right]ij}^{h}\}_{i=1}^{n(\ell)}$ across all available views. View-specific dimensionality $d_{\left[\ell\right]}^{h}$ and target cluster counts $c(\ell)$ are established during this phase, enabling a seamless transition to the initialization and iterative learning phases. Algorithm 2 provides the complete implementation framework for this comprehensive data preparation protocol.
@@ 4.6.2 Initialization Procedures
  • Initialize global view weights $V_{global}^{(0)}$ with an equal distribution across views.
@@
For detailed implementation, refer to Algorithm 3. This algorithm outlines the server- and client-side initialization procedures, ensuring that all clients are prepared for the federated learning process.
@@ 4.6.3 Main Federated Learning Process
The main federated learning process involves multiple communication rounds between the server and the clients. It ensures that each client updates its local model based on the global model parameters received from the server. The clients perform local computations, including heat-kernel enhanced clustering, and send their model updates back to the server. The server aggregates these updates to refine the global model, which is then redistributed to all clients for the next round of learning. The main federated learning process is structured as follows:
@@
  • The process repeats for a predefined number of communication rounds or until the convergence criteria are met.
In this framework, two servers are used to coordinate the federated learning process and manage the global model updates. The first server collects model updates from all clients, aggregates them, and updates the global cluster centers and view weights. The second server is responsible for distributing the updated global model parameters back to the clients for the next round of learning. The unified architecture of the two servers allows for efficient coordination and communication throughout the federated learning process. As a final step, the server evaluates the global model’s performance and adjusts the training strategy as needed. As it reaches convergence, the server outputs the final global model and personalized models for each client, along with their final membership matrices. In conclusion, the user has the capability of implementing a pattern of federated learning across a variety of client devices within a production environment. This ensures that the models are trained and updated without compromising data privacy. A collective, unifying framework is established to facilitate the seamless integration and deployment of federated learning models. This framework comprehensively addresses the lifecycle of federated learning, encompassing model training and deployment. It facilitates efficient collaboration among clients by providing the necessary tools and protocols.
@@
  • Server-Side Components: The servers manage the global model, coordinate communication between clients, and handle the aggregation of model updates.
@@ 4.6.4 Server Aggregation and Distribution
For the server to efficiently aggregate and distribute model updates, the following steps are implemented:
@@
The aggregation process ensures that the global model is a representative update based on the contributions of all clients. This is crucial for maintaining the performance and accuracy of the federated learning system. For server aggregation, we utilize a weighted averaging approach, where the contributions from each client are weighted based on their local data size or importance. The mathematical formulation for the weighted aggregation can be expressed as follows:
@@ 4.6.6 Statistics Aggregation
The statistical analysis also includes the computation of weighted cluster centers and view weights for each client:
@@
These statistics provide insights into the local data distribution, cluster quality, and the importance of views for each client. They can be used to inform the server about the clients’ data characteristics and model performance. These statistics are shared with the server, which can use them to adjust aggregation weights, normalize global parameters, or detect data heterogeneity. For example, the server may assign higher aggregation weights to clients with lower variance or use the global mean and variance to standardize cluster centers and view weights:
@@
where $\mu_{global}^{h}$ and $\sigma_{global}^{h}$ are the global mean and variance for view $h$. These global statistics can be used to normalize the local model updates, ensuring that the aggregated model is robust to variations in client data distributions. The server can then use these statistics to adjust the aggregation process, ensuring that the global model is robust and representative of the clients’ data. This statistical aggregation can also help identify outliers or anomalies in the client updates, allowing for more informed decision-making during the aggregation process.
@@ 4.6.8 The Computational Complexity of FedHK-MVFC
$O\left(\sum_{\ell=1}^{M}\left[n(\ell)^{2}d_{\left[\ell\right]}^{h}s(\ell)+n(\ell)c(\ell)s(\ell)+n(\ell)d_{\left[\ell\right]}^{h}c(\ell)s(\ell)\right]\right)$ (87)
@@ 5.1 Implementation of Privacy-Preserving Mechanisms
5.1.1 Privacy-Preserving Mechanisms in FedHK-MVFC
The FedHK-MVFC framework implements a comprehensive privacy-preserving strategy that combines three complementary mechanisms: data locality, differential privacy, and secure aggregation. These mechanisms function in a coordinated manner to ensure the security of sensitive patient information throughout the federated learning process while facilitating effective collaborative clustering.
Data Locality and Federated Model Updates. In FedHK-MVFC, raw patient data never leaves the local hospital premises. Each hospital $\ell$ maintains its local multi-view dataset $X_{[\ell]}$ and only shares aggregated model parameters. The privacy-preserving update mechanism can be formalized as:
@@ 5.1.1 Privacy-Preserving Mechanisms in FedHK-MVFC
where no raw data $x_{[\ell]ij}^{h}$ is transmitted. For example, Hospital A with ECG data, Hospital B with MRI scans, and Hospital C with genetic profiles only exchange aggregated cluster centers and view weights, not individual patient records. This fundamental principle ensures compliance with healthcare privacy regulations such as HIPAA and GDPR.
Differential Privacy Integration. To provide formal privacy guarantees beyond data locality, FedHK-MVFC incorporates differential privacy by adding calibrated noise to model updates before transmission. The differentially private cluster center update becomes:
@@
This mechanism ensures that individual patient contributions cannot be inferred from the shared model updates, providing mathematically provable privacy guarantees under the $(\epsilon,\delta)$-differential privacy framework.
Secure Aggregation Protocol. Beyond differential privacy, FedHK-MVFC employs a secure aggregation protocol where individual client contributions are cryptographically protected during the server-side aggregation process. The server computes:
@@
where $\text{Encrypt}(\cdot)$ represents homomorphic encryption and $\text{SecAgg}(\cdot)$ denotes the secure aggregation function that allows computation on encrypted data without decryption. This cryptographic layer ensures that even if the communication channel or the aggregation server is compromised, individual hospital contributions remain protected. The combination of these three mechanisms—data locality, differential privacy, and secure aggregation—provides defense-in-depth for privacy preservation, ensuring that FedHK-MVFC meets the stringent requirements of privacy-sensitive healthcare applications.
5.1.2 FedHK-MVFC Flowchart
To provide a comprehensive visual representation of our proposed algorithms, we present a flowchart illustrating the FedHK-MVFC (Algorithm 4) method. Figure 2 illustrates the federated learning workflow, showing the interaction between the central server and distributed clients throughout multiple communication rounds.
Start: Initialize Global & Local ModelsServer: Broadcast $V_{global}^{(t)}$, $A_{global}^{(t)}$Clients: Personalized Init via Eqs. 78, 79Local Updates ($e=1$ to $E$)Compute FedH-KC: $\delta_{[\ell]ij}^{h}$ (Eq. 37)Update $U_{\ell}$, $A_{\ell}$, $V_{\ell}$ (Eqs. 43-45)Compute Statistics: $S_{\ell}^{quality}$, $S_{\ell}^{centers}$, $S_{\ell}^{views}$Send Statistics to $Server$: Aggregate via Eqs. 74, 75Update Personalization: $\gamma_{\ell}$, $\rho_{\ell}$Converged or t=T1t=T-$1t\leftarrow t+$1Output: Global & Personalized ModelsYesNo
Figure 2: Flowchart of the FedHK-MVFC Algorithm (Algorithm 4). The federated workflow involves iterative server-client communication with local heat kernel-enhanced clustering and global model aggregation.
These flowcharts provide visual representations of the algorithmic workflows, highlighting the key differences between centralized HK-MVFC and federated FedHK-MVFC approaches. The centralized version focuses on direct optimization of clustering parameters, while the federated variant incorporates server-client communication, personalization mechanisms, and privacy-preserving aggregation strategies.
5.1.3 Privacy-Preserving Mechanisms in Practice
To demonstrate the practical implementation of privacy-preserving mechanisms in FedHK-MVFC, we present a concrete medical federated scenario involving three hospitals collaborating on cardiovascular patient clustering while maintaining strict privacy compliance.
Scenario Configuration: Consider three hospitals with heterogeneous patient data:
  • Hospital A: 500 patients with electrocardiogram (ECG) records (12 leads each)
  • Hospital B: 300 patients with magnetic resonance imaging (MRI) data (1024 voxel features)
  • Hospital C: 400 patients with genetic profiles (50 SNP markers)
Privacy Implementation: FedHK-MVFC employs differential privacy with privacy budget $\epsilon=1.0$. Each hospital computes local heat-kernel coefficients and cluster updates, adding calibrated Gaussian noise with variance:
$\sigma^{2}=\frac{2\log(1.25/\delta)}{\epsilon^{2}}=\frac{2\log(1.25/10^{-5})}{1.0^{2}}\approx 23.03$ (92)
Encrypted model parameters are then transmitted to the central server, which performs secure aggregation without accessing individual hospital data. Updated global parameters are redistributed for the subsequent federated round.
Adaptive Privacy Budget Allocation: To balance privacy protection with model utility across communication rounds, FedHK-MVFC implements adaptive privacy budget allocation. The total privacy budget $\epsilon_{total}$ is distributed as:
@@ 5.1.3 Privacy-Preserving Mechanisms in Practice
where $T$ denotes the total number of communication rounds and $t$ is the current round. This allocation strategy provides stronger privacy guarantees in early rounds when model updates contain more informative signals, while maintaining convergence properties throughout the training process.
Communication Efficiency and Privacy Trade-offs: The privacy-preserving mechanisms in FedHK-MVFC achieve a favorable balance between privacy protection, communication efficiency, and clustering accuracy. Compared to traditional federated approaches, FedHK-MVFC reduces communication overhead by 70% through:
@@
Privacy-utility analysis demonstrates that with $\epsilon=1.0$, clustering accuracy decreases by only 2-3% compared to the non-private centralized version, while providing formal privacy guarantees compliant with healthcare regulations such as HIPAA. This minimal performance degradation validates the effectiveness of heat-kernel enhanced clustering in preserving geometric structure even under differential privacy constraints.
@@ 6.1 Synthetic Multi-View Data Generation Framework
The synthetic data generator produces two-view, four-cluster datasets, with each view exhibiting distinct geometric patterns. This design philosophy dictates that the clustering algorithms must effectively integrate complementary information across views, rather than relying on a single dominant view. In order to achieve this objective, view 1 is generated based on the principles of parametric geometric structures. As indicated by View 2, complementary geometric patterns emerge, thereby posing a challenge to the clustering process.
6.1.1 Mathematical Formulation of View 1
The initial configuration, designated as ”View 1,” comprises four distinct clusters, each characterized by a unique geometric configuration and distribution. Cluster 1 assumes a circular form, Cluster 2 adopts an elongated horizontal ellipse configuration, Cluster 3 manifests as a crescent shape, and Cluster 4 displays an S-curve or spiral pattern. The generation of these four patterns is achieved through the implementation of parametric equations, which serve to define the spatial distribution of data points within each cluster. This concept is synthetically illustrated through the real-world patient records stored by one hospital. The mathematical formulations employed to generate each cluster are as follows: